A Diesel cycle operates with a compression ratio of 13.5 and with a cutoff

occurring at 6% of the stroke. State 1 is defined by 14 psia and 140 F. For the hot air

standard with k=1.34 and for an initial volume of 1 cu.ft. compute a) Qr, BTU, b)

Wnet, BTU, c) thermal efficiency, %, d) mean effective pressure, psi. e) For a rate of

circulation of 1000 cubic feet per minute, compute the horsepower


1
Expert's answer
2022-03-10T18:01:09-0500

Solution;

Given;

"r=13.5"

k=1.35

"T_1=140+460=600\u00b0R"

"V_1=1cuft"

"C_v=\frac{R}{k-1}=\frac{53.34}{778(1.34)}=0.2016Btu/lb°R"

"C_p=kC_v=1.34\u00d70.2016=0.3602Btu\/lb\u00b0R"

"P_2=P_1(r)^{k-1}=14(13.5)^{0.34}=33.833psia"

"P_2=P_3"

"V_2=\frac{V_1}{13.5}=\frac{1}{13.5}=0.074ft^3"

T"T_2=T_1r^{k-1}=600\u00d713.5^{0.34}=1454\u00b0R"

"V_3=V_2+0.06V_D"

"V_3=0.074+0.06(1-0.074)=0.1298ft^3"

"T_3=T_2(\frac{V_3}{V_2})^{k-1}"

"T_3=1454(\frac{0.1297}{0.0741})^{0.34}"

"T_3=2445\u00b0R"

"T_4=T_3(\frac{V_3}{V_2})^{k-1}=2545(\frac{0.1297}{1})^{0.34}"

"T_4=1271\u00b0R"

"P_4=P_3(\frac{V_3}{V_4})^k"

"P_4=457.9(\frac{0.1297}{1})^{1.34}"

"P_4=29.7psia"

"Q_r=mC_v(T_1-T_4)"

"Q_r=0.063\u00d70.2016(600-1271)"

"Q_r=8.52Btu"

"W_{net}=Q_a-Q_r"

"Q_a=mC_p(T_3-T_2)"

"Q_a=0.063\u00d70.2702(2545-1454)=18.57Btu"

"W_{net}=18.57-8.52=10.05Btu"

"\eta=\frac{W}{Q_a}=\frac{10.05}{18.57}=0.5412""P_m=\frac{W}{V_1-V_2}=\frac{10.05×778}{(1-0.0741)×144}"

"P_m=58.64psi"

"W_r=\frac{10.05×1000}{42.4}=237hp"